OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 18, and minor axis of length 10.

OpenStudy (anonymous):

I really need help, im in precalculus and my homework problems are hard for me,

OpenStudy (mertsj):

Where is the center of this ellipse?

OpenStudy (anonymous):

(0,0)

OpenStudy (mertsj):

The major axis is 18. The major axis is 2a. Do you know that?

OpenStudy (anonymous):

yes i got that, i know its either x^2/5 +y^2/9=1 or x^2/9 +y^2/5=1 im just getting the order so confused

OpenStudy (mertsj):

The minor axis is 10. The minor axis is 2b. Do you know that?

OpenStudy (mertsj):

It would be very helpful if you would just answer my questions. Do you know that the major axis is 2a or not?

OpenStudy (anonymous):

yes i get that:), so im confused because i dont know which equation i would use for this , x^2/a^2 +y^2/b^2 or the a and b switched

OpenStudy (mertsj):

\[\frac{x^2}{25}+\frac{y^2}{81}=1\]

OpenStudy (anonymous):

ohhhhhh! your so smart!, i was thinking i would squarerroot the bottoms of both of those